OpenStudy (anonymous):

You have the numbers 1–24 written on slips of paper. If you choose one slip at random, what is the probability that you will not select a number which is divisible by 3?

OpenStudy (anonymous):

very similar to the last one divisible by three would be 3,6,9,12,15,18,21,24 and there are 8 of them divisible by 3 and so 24 - 8 = 16 not divisible by 3 so \(\frac{16}{24}=\frac{2}{3}\)